# V-FLUTE: Visual Figurative Language Understanding with Textual Explanations

### v1

@@ 4 Experiments
We empirically study how several baseline models perform on the task of explainable visual entailment. We investigate both off-the-shelf and fine-tuned model performance.
@@ 4.1 Models
We select a variety of models for our study (see taxonomy in Appendix, Figure 10). For off-the-shelf models, we explore both open and API-based models. For open models, we select the (current) state-of-the-art LLaVA-1.6 models Liu et al. (2024). LLaVA is one of the simplest, yet one of the most high-performing VLM architectures currently available. It utilizes a pretrained large language model (e.g., Mistral-7B Jiang et al. (2023)) and a vision-language cross-modal connector (e.g., an MLP layer) to align the vision encoder (e.g., CLIP Radford et al. (2021)) outputs to the language models. We select LLaVA-1.6 models in their 7B and 34B configurations (LLaVA-v1.6-7B and LLaVA-v1.6-34B respectively) and refer to them as LLaVA-ZS-7B and LLaVA-ZS-34B. Both models have been instruction-tuned on less than 1M visual instruction tuning samples to act as general language and vision assistants. It should, however, be noted that these models do not currently support few-shot multimodal prompting.
In addition to zero-shot testing, we also test these models using Compositional Chain-of-Thought Prompting proposed by Mitra et al. (2023). The method first prompts the model to generate a scene graph and then utilizes that scene graph in another prompt to answer the relevant question. The method works zero-shot without requiring fine-tuning. We refer to these models as LLaVA-ZS-7B-SG and LLaVA-ZS-34B-SG for the 7B and 34B LLaVA configurations described above.
For API-based models, we select three widely available state-of-the-art VLMs: Claude-3 Opus (claude-3-opus-20240229)Anthropic (2024), GPT-4 (gpt-4-1106-vision-preview) OpenAI (2023) and GeminiPro (gemini-pro-vision)Team et al. (2023). We refer to GPT-4 as the “teacher” model as most candidate explanations were generated with it.
For fine-tuned models, we focus on fine-tuning LLaVA-1.5-7B model Liu et al. (2023a) (the fine-tuning code for 1.6 model is not available during the time the paper was written). To minimize bias for a single instruction, we fine-tune and evaluate the models on a set of 21 instruction paraphrases (see Appendix Table 8). Three model configurations are tested:
LLaVA-eViL is a checkpoint of LLaVA-v1.5-7B further fine-tuned on the eViL (e-SNLI-VE) dataset for explainable visual entailment Kayser et al. (2021) converted to the instruction format. We removed neutral label instances, which resulted in 275,815 training instances and 10,897 validation instances.
@@
LLaVA-eViL+VF is the same checkpoint fine-tuned on both eViL and V-FLUTE.
All hyperparameters are in Appendix C.
@@ 4.2 Automatic Metrics
Similarly to prior work Chakrabarty et al. (2022) we utilize both classic F1 score and an adjusted score that accounts for explanation quality: F1@ExplanationScore. The ExplanationScore computes the average between BERTScore Zhang* et al. (2020) based on the microsoft-deberta-xlarge-mnli model He et al. (2021); Williams et al. (2018) and BLEURT Sellam et al. (2020) based on the BLEURT-20 Pu et al. (2021). Since our goal is to ensure models provide an answer for the right reasons, ideally, we would only count predictions as correct when the explanation is also correct. Hence, we report F1@0 (simply F1 score), F1@53 (only predictions with explanation score 53 are considered correct), and F1@60. Thresholds are selected based on human evaluation of explanation quality in Section 5.3.
@@ 4.3 Automatic Evaluation Results
Table 3 shows the results based on the automatic evaluation. We also include results per phenomenon in Appendix F and the drop in performance when accounting for explanations score in Figure 6. Our results inform the following insights:
@@
Fine-tuning on V-FLUTE leads to best classification performance on average across datasets.
Our strongest fine-tuned model (LLaVA-7B-eViL+VF) outperforms the best off-the-shelf model (GPT-4-5shot) in terms of the F1@0 score (; all values reported via paired bootstrap test Koehn (2004)), and performs competitively when incorporating the explanations quality with GPT-4 leading slightly (F1@60 of 49.81 vs. 48.80 for the best fine-tuned model), which is expected as GPT-4 is the teacher model with which the majority of the explanation candidates were generated. Adding the e-ViL dataset improves the performance slightly compared to only fine-tuning on V-FLUTE. Fine-tuning merely on e-ViL improves over a random baseline; however, the explanations are of poor quality.
We also utilize a hypothesis-only baseline Poliak et al. (2018) by including a model fine-tuned on the V-FLUTE dataset, but without the relevant image (with a white square as an input instead, denoted as Image). Fine-tuning on the full V-FLUTE dataset shows an improvement of over 8 points in F1@0 (better with ), suggesting VLMs benefit from visual information when dealing with figurative phenomena and do not just rely on the input text to make their prediction.
Open zero-shot instruction-tuned models are lagging behind API-based models, but scene graph prompting improves performance.
LLaVA-7B and 34B lag behind Claude 3 and GPT-4 in zero-shot settings. However, scene graph prompting improves the zero-shot performance of the LLaVA-based models, allowing them catch up to zero-shot API model performance (Gemini and Claude 3). The explanations generated by these models tend to overly focus on the contents of the scene graph rather than the underlying figurative phenomena, possibly causing a decrease in explanation score (and consequently in F1@60). The few-shot API models outperform zero-shot API models, and are better than all configurations of open models in F1@0, 53, 60, indicating the effectiveness of few-shot prompting (not available for LLaVA-based models as of now).
Performance for models decreases when taking into account explanation quality.
We plot the relative percentage decrease between F1@0 and F1@60 for LLaVA-eViL-VF, LLaVA-34B-SG, and GPT-4-5shot in Figure 6. Higher relative drop indicates higher difficulty of generating the correct explanation. For all models, we see a substantial decrease in performance, especially on challenging phenomena such as Humor (NYCartoons). For Metaphor (IRFL), Humor (MemeCap) and Idiom (IRFL) subsets GPT-4 exhibits the lowest relative performance drop, while for Metaphor (HAIVMet), Humor (NYCartoons) and Sarcasm (MuSE) the fine-tuned model has the lowest drop.
We can see that the percentage drop is substantially higher for all models for the HAIVMet subset rather than the IRFL dataset, which contains metaphors in the image rather than in the text. This suggests it is harder for models to generate correct explanations when the figurative meaning is contained in the image rather than in the text, indicating the need to expand current datasets to include images with figurative meaning.
4.4 Human Baseline
Phenomenon | Dataset |
Metaphor /Similes | HAIVMET | 78.84 | 81.25 | ||||
94.36 | 77.78 | |||||
Idioms |
89.26 | 49.74 | ||||
Sarcasm | MuSE | 68.89 | 85.42 | ||||
Humor | MemeCap | 100.0 | 78.03 | ||||
NYCartoons | 71.43 | 47.83 | |||||
Overall | 89.09 | 77.26 |
@@ 5 Human Evaluation and Error Analysis
We conduct human evaluation of generated explanation to more reliably assess their quality and identify key errors in multimodal figurative language understanding. We recruit two expert annotators with background in linguistics for the task and sample 95 random instances from the test set. For each instance, we first provide the annotators with the image, claim and reference explanation and ask the annotators to choose the right label. If the annotator succeeds, they can view the rest of the task, which consists of 3 explanations from our top models by F1@0 in each category: LLaVA-eViL-VF, LLaVA-34B-SG, GPT-4-5shot. The explanations are taken for both correct and incorrect model predictions. For each explanation, we ask whether the explanation is adequate (accurate, correct, complete and concise). If not, we ask them to identify one of the three main types of errors based on the following taxonomy:
Hallucination: explanation is not faithful to the image, indicating difficulties with basic visual comprehension (see prediction of a blunt tip when the pencil tip is actually sharp in row 1 of Table 5).
Unsound reasoning: sentences do not adhere to natural logic or violate common sense (e.g., concluding than an upwards arrow and lots of money imply an economic crisis, see row 3).
Incomplete reasoning: while overall the explanation makes sense, it does not address the key property reasons why the image entails or contradicts the claim (for example, does not address the figurative part in the image, see row 2).
Too Verbose: the explanation is too verbose to the point it would interfere rather than help one decide the correct label.
@@ 5.1 How Do Models Perform According to Humans?
In Table 6, we show adequacy and preference rates for explanations from the 3 systems, where an explanation is deemed adequate if both annotators agreed it is, and inadequate if both agreed it is not. The preference percentage is also taken among instances where the annotators agreed that the model’s explanation is preferred among all the adequate explanations. The average IAA using Cohen’s is 0.47, indicating moderate agreement Cohen (1960). We observe that the teacher model is leading in terms of the adequacy of the explanations and preference rate, as expected from a larger system equipped for higher quality reasoning and generation capabilities. Yet still only half of its explanations are considered adequate. This further confirms that despite impressive performance on the F1@0 scores, the models are not yet capable of producing adequate textual explanations in many instances.
@@ 5.2 What Errors Do Models Make?
We also analyze to understand what type of errors do each model make when they are considered not adequate in the above evaluation. In Figure 7, we illustrate the normalized frequency of error types when both annotators agree that the explanation is not adequate (i.e., out of all errors for this model, what percentage is each type of error?). In general, annotators did not consider verbosity to be a major issue of the systems. For GPT-4, the leading error type is hallucination, indicating the need to improve faithful image recognition even in the most advanced models. For the fine-tuned model and LLaVA-34B-SG, the main error type is unsound reasoning, indicating that it is challenging for the models to reason about multimodal figurative inputs consistently.
5.3 How Well Does the Explanation Score Predict Human Judgment on Adequacy?
We explore whether the proposed explanation score can capture human judgement of explanation adequacy. We collect all instances where both annotators agreed on the adequacy judgement for the explanation. We evaluate if the explanation score described in Section 4.2 can act as a good predictor for the human adequacy judgment. We find that the area under the Precision-Recall curve is 0.79, and the maximum F1 score is 0.77, obtainable at the explanation score threshold of 0.53. Hence, we use this threshold to report the results in Table 3. We also use the threshold of 0.6 since it maximizes F1 such that both precision and recall are above 0.75.
@@ 6 Conclusions
We introduce a high-quality dataset for understanding figurative phenomena in multimodal input, V-FLUTE, framed as an explainable visual entailment. Our dataset consists of 6,027 image, claim, label, explanation instances spanning a variety of figurative phenomena such as metaphor, idiom, simile, sarcasm, and humor. We use this dataset to benchmark the performance of state-of-the-art vision-language models using both automatic and human evaluation and to identify critical areas of improvement for VLMs for this task.
Appendix A Dataset Statistics
@@
Average length of a claim in V-FLUTE is 61 characters. Average length of an explanation is 367 characters. Figure 8 shows the distribution of claim lengths, and Figure 9 shows the distribution of explanation lengths by source dataset. We manually verified that the outlier instances are correct.
Appendix B API models Hyperparameters
B.1 Claude
@@ B.3 Gemini
Appendix C Fine-tuning Hyperparameters
We utilize LoRA Hu et al. (2022) to fine-tune the models. We utilize the same hyperparameters for all fine-tunes outlined in Appendix C and use early stopping based on a V-FLUTE validation set to prevent overfitting. Due to size of the data, we only train for 3 epochs in both eViL and eVIL+VFLUTE variations, but we saved intermediate steps and chose the checkpoint where the validation loss is lowest.
@@ Appendix C Fine-tuning Hyperparameters
Appendix D Prompts for LLMs
D.1 HAIVMET
D.1.1 One-shot Prompt for generating explanations

### v2

@@ 4 Experiments
We empirically study how several baseline models perform on the task of explainable visual entailment. We investigate both off-the-shelf and fine-tuned model performance. We provide human baseline performance in Appendix K. Hyperparameters are provided in Appendix E.
@@ 4.1 Models
We select a variety of models for our study (see taxonomy in Appendix, Figure 10). For off-the-shelf models, we explore both open and API-based models. For open models, we select the (current) state-of-the-art LLaVA-1.6 models Liu et al. (2024). LLaVA is one of the simplest, yet one of the most high-performing VLM architectures currently available. It utilizes a pretrained large language model (e.g., Mistral-7B Jiang et al. (2023)) and a vision-language cross-modal connector (e.g., an MLP layer) to align the vision encoder (e.g., CLIP Radford et al. (2021)) outputs to the language models. We select LLaVA-1.6 models in their 7B and 34B configurations (LLaVA-v1.6-7B and LLaVA-v1.6-34B respectively) and refer to them as LLaVA-ZS-7B and LLaVA-ZS-34B. Both models have been instruction-tuned on less than 1M visual instruction tuning samples to act as general language and vision assistants. We also utilize Compositional Chain-of-Thought Prompting proposed by Mitra et al. (2023) denoted by LLaVA-ZS-7B-SG and LLaVA-ZS-34B-SG (see description and results discussion in Appendix H).
For API-based models, we select three widely available state-of-the-art VLMs: Claude-3 Opus (claude-3-opus-20240229)Anthropic (2024), GPT-4 (gpt-4-1106-vision-preview) OpenAI (2023) and GeminiPro (gemini-pro-vision)Team (2023).
For fine-tuned models, we focus on fine-tuning the LLaVA-1.5-7B model222Fine-tuning code for 1.6 model was not published as of writing of this paper. Liu et al. (2023a). To minimize bias for a single instruction, we fine-tune and evaluate the models on a set of 21 instruction paraphrases (see Appendix Table 7). Three model configurations are tested:
@@
LLaVA-eViL and LLaVA-eViL+VF are checkpoints of LLaVA-v1.5-7B further fine-tuned on the eViL (e-SNLI-VE) dataset for explainable visual entailment Kayser et al. (2021) converted to the instruction format or on both eViL and V-FLUTE. We removed neutral label instances, which resulted in 275,815 training instances and 10,897 validation instances.
@@ 4.2 Automatic Metrics
Since our goal is to ensure models provide an answer for the right reasons, ideally, we would only count predictions as correct when the explanation is also correct. Based on prior work Chakrabarty et al. (2022), we use both the standard F1 score and an adjusted score that accounts for explanation quality: F1@ExplanationScore. The ExplanationScore computes the average of BERTScore Zhang* et al. (2020) and BLEURT Sellam et al. (2020) between model-generated and reference (V-FLUTE) explanations. We report F1@0 (simply F1 score), F1@53333Thresholds selected based on human evaluation of explanation quality in Appendix L. (all predictions with ExplanationScore 53 are considered incorrect) and F1@60.
@@ 4.3 Automatic Evaluation Results
We include results per phenomenon in Appendix J, discussion on CoT prompting in Appendix H and additional models in Appendix I. Table 3 shows the results, informing the following insights:
@@
A literal visual entailment dataset does not solve the figurative visual entailment task.
Fine-tuning only on e-ViL barely improves over a random baseline (54.34 F1@0) and underperforms compared with the models fine-tuned on V-FLUTE (72.78 F1@0). Moreover, the explanations are of poor quality (0.55 F1@60). This indicates that models trained on a literal visual entailment task struggle to generalize to figurative meaning, supporting the challenging nature of our dataset.
The strongest model fine-tuned on V-FLUTE (LLaVA-7B-eViL+VF) outperforms the best off-the-shelf model (GPT-4-5shot) in terms of the F1@0 score (444 values reported via paired bootstrap test Koehn (2004)). It performs competitively when incorporating the reference-based ExplanationScore, with GPT-4 leading slightly as it is the model with which the candidate explanations were generated.
When figurative meaning is in the image rather than text, models perform worse.
We plot the relative percentage decrease between F1@0 and F1@60 for LLaVA-eViL-VF, LLaVA-34B-SG, and GPT-4-5shot in Figure 5. Higher performance drop indicates higher difficulty of generating the correct explanation. For all models, we see a substantial decrease in performance, especially on challenging phenomena such as Humor (NYCartoons). The percentage drop is substantially higher for all models for the HAIVMet subset rather than the IRFL dataset, which contains metaphors in the image rather than in the text. This suggests it is harder for models to generate correct explanations when the figurative meaning is contained in the image rather than in the text, indicating the need to expand the presence of figurative phenomena in existing visual datasets.
VLMs benefit from visual information when dealing with figurative phenomena and do not just rely on the input text to make their prediction.
We utilize a hypothesis-only baseline Poliak et al. (2018) by including a model fine-tuned on the V-FLUTE dataset, but with a white square as the image input, denoted as Image. Fine-tuning on the full V-FLUTE dataset shows an improvement of over 8 points in F1@0 (better with ).
@@ 5 Human Evaluation and Error Analysis
We conduct human evaluation of generated explanations to assess their quality and identify key errors in reasoning about multimodal figurative meaning. We recruit two expert annotators with background in linguistics for the task and sample 95 random instances from the test set. For each instance, we first provide the annotators with the image, caption and reference explanation and ask the annotators to choose the right label. If the annotator succeeds, they can view the rest of the task, which consists of 3 explanations from our top models by F1@0 in each category: LLaVA-eViL-VF, LLaVA-34B-SG, GPT-4-5shot. The explanations are taken for both correct and incorrect model predictions. For each explanation, we ask whether the explanation is adequate (accurate, correct, complete and concise). If not, we ask them to identify one of the errors based on the following taxonomy:
Hallucination: explanation is not faithful to the image, indicating difficulties with visual comprehension (e.g., generates “blunt tip” when the pencil tip is actually sharp in row 1 of Table 4).
Unsound reasoning: sentences do not adhere to natural logic or violate common sense (e.g., concluding that an upwards arrow and lots of money imply an economic crisis, see row 3).
Incomplete reasoning: while overall the explanation makes sense, it does not address the key property reasons why the image entails or contradicts the caption (for example, does not address the figurative part in the image, see row 2).
Verbosity: the explanation is too verbose.
@@ 5.1 How Do Models Perform According to Humans?
In Table 5, we show adequacy and preference rates for explanations from the 3 systems, where an explanation is deemed adequate or preferred if both annotators agreed it is, and inadequate if both agreed it is not. The average IAA using Cohen’s is 0.47, indicating moderate agreement Cohen (1960). We observe that the teacher GPT-4 model is leading in terms of the adequacy of the explanations and preference rate, as expected from a larger system. Yet still only half of its explanations are considered adequate, confirming that despite good performance on the F1@0 scores, the models are not yet capable of producing adequate textual explanations in many instances. 555Note that during the human-AI collaborative dataset creation 1) the LLM is conditioned on the correct label, 2) its explanation is edited by an expert annotator.
@@ 5.2 What Errors Do Models Make?
We perform an analysis of the types of errors from each model when the explanations are considered inadequate in the above evaluation. In Figure 6, we illustrate the normalized frequency of error types when both annotators agree that the explanation is not adequate (i.e., out of all errors for this model, what percentage is each type of error?). Overall, the annotators did not consider verbosity to be a major issue of the systems. For GPT-4, the leading error type is hallucination, indicating the need to improve faithful image recognition even in the most advanced models. Comparing LLaVA-34B-SG and the fine-tuned model, we see that for the scene graph model a larger percentage of errors is due to incomplete reasoning (possibly due to focusing on the scene graph description rather than the underlying figurative phenomena). For both models, the main error type is unsound reasoning, indicating difficulty for the models to consistently reason about multimodal figurative inputs.
@@ 6 Conclusions
We introduce a novel dataset for understanding figurative meaning in multimodal input, V-FLUTE, via an explainable visual entailment task. Our dataset consists of 6,027 image, caption, label, explanation instances covering diverse phenomena. We find that VLMs struggle to generalize from literal to figurative meaning, particularly in images, and identify error types in VLM reasoning about multimodal figurative phenomena.
7 Ethics
Following prior work in human-AI collaboration for complex text and image generation Chakrabarty et al. (2022, 2023); CH-Wang et al. (2023); Saakyan and Muresan (2023), we opt for an expert-AI collaboration framework where experts edit the initial generations by the language model. Expert feedback is essential to improve the quality of the data, as previous work has identified that crowdworkers on platforms such as Amazon Mechanical Turk could be unreliable for open-ended generation tasks Karpinska et al. (2021), and might even rely on ChatGPT to provide their answers Veselovsky et al. (2023). To mitigate these effects, in this work, annotators were recruited through the Upwork platform, allowing to select for relevant level of expertise and verify, e.g., educational and professional background of the annotators. All recruited annotators have significant background in figurative language understanding and have formal educational background in linguistics or literature. All of the annotators are fluent or native/bilingual level in English. Workers on UpWork were informed that that the work they were doing was going to be used for research purposes. All are fairly compensated with USD $20 to $25 per hour with self-reported time needed to complete the tasks. The total budget for the annotation and GPT-4 generations was USD. We estimate that it would take approximately 3 times longer to complete the annotation task without the pre-generated explanation, so we estimate that the cost would have at least tripled if the human-AI collaboration approach was not utilized. Workers were paid their wages in full immediately upon the completion of their work. All data collected by human respondents were fully anonymized. We do not report demographic or geographic information, given the limited number of respondents, so as to maintain full anonymity.
8 Limitations
We would like to acknowledge the following limitations of our work. The textual explanations in V-FLUTE dataset were generated with the help of the strongest LLM available at the time of writing the paper, GPT-4. Despite our best efforts in mitigating biases with expert human verification, idiosyncrasies pertaining to GPT-4 outputs may still be present in the text. This means that it is potentially possible for the underlying biases of source datasets of language model generations to propagate into our resource, which we wish to mitigate by carefully examining each dataset instance by one of the 3 expert annotators.
Reference-based evaluation has fundamental flaws, such as not considering all possible explanations, which would be impossible to collect. However, current reference-free metrics for free-text rationales may still have flaws such as bias toward length or the evaluator LLM Stureborg et al. (2024); Raina et al. (2024); Huang et al. (2024); Chiang and Lee (2023); Wei et al. (2024). When evaluating textual explanations against these references, as is the case with any reference-based evaluation, there may also be a preference towards models which output text closer in distribution to the GPT-4 model. Because of that, it is important to utilize the data set in order to compare models other than the teacher model and pay more attention to the F1@0 scores, which represent simple classification scores and do not require the outputs to be similar in distribution. In terms of pure F1 score performance, GPT-4 underperforms the fine-tuned model, and performs very closely with Gemini and Claude that were not used to generate the data, with less than 2% difference (see column F1@0, Table 4). Although we showed a relatively high predictive power of automatic explanation scores to predict human judgements (see Appendix L), future work may focus on increasing reliability of reference-based and reference-free textual explanation evaluation methods.
We also note that the images from the HAIVMet dataset Chakrabarty et al. (2023) are AI generated. However, the majority of the remaining images in V-FLUTE are not AI generated but are naturally occurring or created by humans. However, to mitigate potential biases from AI-generated images, all instances of the data were examined during the expert verification stage, as described in the article.
Label predictions by language models can vary significantly with slight differences in prompt wording Sclar et al. (2023), which is why during fine-tuning and inference we utilize over 20+ different templates of instructions (see Table 7). Nevertheless, it is important to consider the models’ explanations to better assess their understanding of the phenomena, which we hope to enable with our explainable figurative visual entailment dataset.
Appendix A Dataset Pipeline Illustrations
We illustrate the pipelines for creating V-FLUTE from each underlying figurative phenomenon in Figures 7, 2, 3, 4.
Appendix B Details on Expert Verification
We follow the same procedure for expert verification of all sub-datasets. We recruit 3 expert annotators with background in figurative language and formal educational background in linguistics or literature on Upwork. We first ask to annotate 10 instances by all 3 annotators to ensure they understand the task. We then ensured a high agreement ( pairwise accuracy) between annotators on a subsample of 100 instances of each dataset, and resolved any disagreements through mutual discussion between the annotators and the authors before proceeding. Finally, each annotator proceeds to annotate roughly of the data.
We provide the annotation interfaces below for HAIVMET (Figure 12), IRFL (Figure 13), MemeCap (Figure 14) and MuSE (Figure 15). In addition, instructions were explained in more detail to the annotators via chat on Upwork (for example, the criteria for correctness and conciseness), and any of their doubts and questions were answered.
Appendix C Dataset Statistics
@@
Average length of a caption in V-FLUTE is 61 characters. Average length of an explanation is 367 characters. Figure 8 shows the distribution of caption lengths, and Figure 9 shows the distribution of explanation lengths by source dataset. We manually verified that the outlier instances are correct.
Appendix D API models Hyperparameters
D.1 Claude
@@ D.3 Gemini
Appendix E Fine-tuning Hyperparameters
LLava-v1.6-6B and 34B respectively utilize instruction-tuned LLMs as their backbone, Mistral-7Binstruct666huggingface.co/mistralai/Mistral-7B-Instruct-v0.1 and Yi-34B777huggingface.co/NousResearch/Nous-Hermes-2-Yi-34B.
We utilize LoRA Hu et al. (2022) to fine-tune the models. We utilize the same hyperparameters for all fine-tunes outlined in Appendix E and use early stopping based on a V-FLUTE validation set to prevent overfitting. For evil and e-ViL+V-FLUTE we only fine-tuned for 2 epochs due to size of the e-ViL dataset and took the best checkpoint based on early stopping on V-FLUTE validation set. For eViL we only fine-tuned for 1 epoch to prevent overfitting. For VFLUTE, we trained for 3 epochs, and VFLUTE -Image for 10 epochs (to ensure performance does not increase even with larger number of epochs), for both we took the best checkpoint based on early stopping.
We utilize 4 NVIDIA A100 40GB GPUs for all experiment.
@@ Appendix E Fine-tuning Hyperparameters
The instructions were almost always followed. If they were not followed during the data creation process, we discarded those instances. For evaluation, we looked at the sample outputs of each model and designed rules to extract the label and the explanation from the output, which was not too difficult since mostly the instructions were followed well. In the rare cases the model failed to follow instructions, that label would likely be incorrect.
@@
Evaluation Hyperparameters
Appendix F Prompts for LLMs
F.1 HAIVMET
F.1.1 One-shot Prompt for generating explanations